OpenStudy (anonymous):

Find the equation of the line that contains the point (5, 8) and has a slope of 1/2. (5, 8) and m=1/2 help?

OpenStudy (anonymous):

Point slope form: Given m and \[\(x_\0\,y_\0\)\] Then the equation of a line is given by: \[y-y_\0\=m(x-x_\0\)\] Plug in and solve for y.

OpenStudy (bahrom7893):

y - y1 = M (x - x1) y - 8 = (1/2) ( x - 5) y = (1/2)x - 5/2 + 8 y = (1/2)x - 1/2

OpenStudy (amistre64):

your latex looks wierd, did it format right on your screen?

OpenStudy (anonymous):

Nah :/ \[(x_0,y_0)\] Then use: \[y-y_0=m(x-x_0)\]

OpenStudy (anonymous):

There we go :)

OpenStudy (amistre64):

thats better :)

OpenStudy (anonymous):

Yeah mine did that too :/ Idk what that was about.

OpenStudy (anonymous):

Ok, apparently that answer is wrong, here is the question word for word "Find the equation of the line containing the give point and slope. (5,8) and m=1/2

OpenStudy (amistre64):

which form of the equation are they after?

OpenStudy (bahrom7893):

hmm that answer is right, maybe you need to pull out 1/2: y = (1/2)( x - 1 )

OpenStudy (amistre64):

y = 1/2 x -1/2(5) +8

OpenStudy (anonymous):

yah its really vague on what they want.

OpenStudy (amistre64):

y = \(\frac{x+11}{2}\) is one way